y''+3y'+2y= 1/(1+ex)


1
Expert's answer
2021-08-05T13:54:14-0400

Related homogeneous differential equation


"y''+3y'+2y=0"

The roots of the characteristic equation are


"r^2+3r+2=0"

"(r+1)(r+2)=0"

"r_1=-2,r_2=-1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-2x}+c_2e^{-x}"

Use the method of variation of parameters


"c_1'e^{-2x}+c_2'e^{-x}=0"

"c_1'(e^{-2x})'+c_2'(e^{-x})'=\dfrac{1}{1+e^x}"


Then


"c_2'e^{-x}=-c_1'e^{-2x}"

"-2c_1'e^{-2x}+c_1'e^{-2x}=\dfrac{1}{1+e^x}"

"c_1'=-\dfrac{e^{2x}}{1+e^x}"

"c_1=-\int\dfrac{e^{2x}}{1+e^x}dx=\ln(e^x+1)-e^x+C_1"

"c_2'=\dfrac{e^{x}}{1+e^x}"

"c_2=\int\dfrac{e^{x}}{1+e^x}dx=\ln(e^x+1)+C_2"

The general solution of the homogeneous differential equation is


"y=e^{-2x}\ln(e^x+1)+C_1e^{-2x}+e^{-x}\ln(e^x+1)+C_3e^{-x}"